Here's another accuplacer practice question that I have no clue about: "write the slope intercept form of the equation (-2,0)and (7,4). The answer that lights up correct is Y=4/9x + 8/9. Can someone tell me why this is correct, and in general , how to write slope intercept forms? Thanks lots.
Answer by london maths tutor(243)

The linear equation form = Y = m X + C
where m = gradient and C = y intercept
gradient = (4-0)/[7-(-2)]
= 4/9
Sub in one of the point (-2,0) and m = 4/9 into the linear equation:
0 = -2(4/9) + C
0 = -8/9 + C
C = 8/9
Therefore, the equation: Y = (4/9) X + (8/9) Hence shown